Operation Summary for January 28, 2011 - January 31, 2011 |
|
The Integrated Luminosity for the period from 1/24/11 to 1/31/11 was 67.3 Pb-1 (inverse picobarns). NuMI reported receiving 8.65E18 protons on target during this same period. Friday Jan 28 The day shift began with Operations monitoring the Tevatron (TEV) store 8458 (75.41E30) and the Recycler stash (310.6E10), with the Antiproton Source (Pbar) stacking (21.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, with the Meson FTBF T-978 (CALICE) experiment off due to a network storm, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, operators began the shift trying to recover from the network storm. Operations terminated store 8458 at 10:40 AM. Cryo system personnel began TEV sector B1 wet engine repairs. Operations prepared the D0 collision hall for an access. After the completion of the wet engine work, Operations put the TEV through a dry squeeze and a wet squeeze. Operations put the TEV into shot setup at 2:27 PM and established store 8461 at 3:51 PM with an average initial luminosity (AIL) of 314.59E30. Operations reported that the Pbar Accumulator (AP50) small chiller developed a leak; operators checked and filled it as necessary throughout the weekend. Operations disabled the NuMI beam twice so that Controls personnel could conduct CAMAC front end repairs to alleviate future network storms. Operations reported sending beam to T-978, NuMI, and MiniBooNE when it was available. During the evening shift, Operations disabled the T-978 beam at 10 PM. Operators reported problems with a Pbar Accumulator PLC (A:ISHUT); Pbar personnel found the PLC code gone and reloaded the program. Saturday Jan 29 The midnight shift began with Operations monitoring store 8461 (110.56E30) and the stash (311.7E10) and the Pbar stack (26.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8461 at 2:41 AM, put the TEV into shot setup, and established store 8462 at 4:02 AM with an AIL of 336.3E30. Operations established beam to T-978 at 4:50 AM. During the day shift, Operations monitored the accelerator complex. During the evening shift, Operations terminated store 8462 at 5:39 PM, put the TEV into shot setup, and established store 8463 at 6:47 PM with an AIL of 327.4E30. Operators tuned Linac and Booster. Operations disabled beam to T-978 when shot setup began. Sunday Jan 30 The midnight shift began with Operations monitoring store 8463 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:26 AM. During the day shift, Operations terminated store 8463 at 9:03 AM, put the TEV into shot setup, and established store 8464 at 10:21 AM with an AIL of 323.03E30. During the evening shift, Operations disabled the T-978 beam at 6 PM. Monday Jan 31 The midnight shift began with Operations monitoring store 8464 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8464 at 12:59 AM, put the TEV into shot setup, and established store 8465 at 2:22 AM with an AIL of 358.5E30. Operations established beam to T-978 at 4:25 AM. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and TEV personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 01/31/2011 email Fermilab |
| Security, Privacy, Legal | |